suppose that the monetary base is 500 billion the reserve deposit ratio is 0.1 and the currency deposit ratio is 0.6 . find the money multiplier and money supply


1
Expert's answer
2021-08-13T10:22:40-0400

Money Multiplier "=" "\frac {CDR+1}{RDR+CDR}" "=" "\frac {0.6+1}{0.1+0.6}" "=" 2.28571


Money supply"=" money multiplier"\u00d7" monetary base

2.28571"\u00d7" 500"=" 1142.8571

Answer"=" 1143